Write a molecular formula for an alkane whose density with respect to hydrogen is 50. (Properly named alkane is evaluated
1
Expert's answer
2020-11-09T14:08:09-0500

"D_{H_2}=\frac{M(C_nH_{2n+2})}{M(H_2)}"

"M(C_nH_{2n+2})=D_{H_2}\times M(H_2)=50\times2=100 \frac{g}{mole}"

"12n+1(2n+2)=100"

"n=7"

Alkane is heptane ("C_7H_{16})" )